How can you solve an equation with one variable and two numbers if they have a negative and positive sign. It will be okay if you give me an example of how to solve an equation. If it`s not a problem, could show step by step please and thank you.
Answer by KneesAndToes(13)

Here is an example:


5x+2 = -13


First you will want to subtract 2 from each side of the equation. 2-2 = 0 and
-13-2 = -15, so


5x = -15


Now you will want to divide both sides by 5 to get the x alone. 5x/5 = x and
-15/5 = -3 so


x = -3 ... 5(-3) = -15 ... -15 = -15


Another example:


4-2x = -2


First subtract 4 from each side. 4-4 = 0 and -2-4 = -6


-2x = -6


Now divide each side by -2. -2/-2 = 1 and -6/-2 = 3. Any variable on it's own can also be thought of with a one in front of it. x=1x or y=1y. If it makes it easier for you, you can put the one in front


1x = 3 or x = 3 ... -2(3) = -6 ... -6 = -6


One more:


10-x = 5


Subtract 10 from each side. 10-10 = 0 and 5-10 = -5.


-x = -5


Now, if you remember that x=1x, you can divide both sides by -1. -1x/-1 = x and
-5/-1 = 5, so


x = 5 ... 10-5 = 5 ... 5 = 5.


Just remember to start by getting the variable alone to one side, and remember the rules of positives and negatives. I hope this was what you meant, and I hope it helped...